OpenStudy (anonymous):

Find the area enclosed by the x-axis and the curve x=t^3, y=1-t^6.

OpenStudy (anonymous):

Since it is just the x-axis... i do not need to subtract anything? just integrate y(dx/dt)dt right?

OpenStudy (anonymous):

this might help http://www.mathwords.com/a/area_under_a_curve.htm

OpenStudy (anonymous):

did you get an answer?

OpenStudy (anonymous):

\[t^3=x\] \[t=\sqrt[3]{x}\] Replace t in the equation for y to get the equation in terms of x \[y=1-(\sqrt[3]{x})^6\] Expand the exponent of 6 to each factor in the expression y=1-x^2 Multiply -1 by the X6^2 inside the parentheses \[y=1-x^2\] Not positive if this is right do you have choices?

OpenStudy (anonymous):

no choices. and I need to use calculus on it. The answer should be a number. to find area I integrate from a to b ... with in this case is 0 to 1 of y (dx/dt) dt. I got 2/3 sq units.. do not know if that is correct.

OpenStudy (anonymous):

@radar can you help?

OpenStudy (anonymous):

thanks for helping me by the way.

OpenStudy (anonymous):

sorry wish I could help your in Geometry

OpenStudy (anonymous):

hehe Multivariate Calculus.

OpenStudy (anonymous):

lol i better not help you :) i might mess you up. Im in geometry just getting into trig

OpenStudy (anonymous):

fun times! Trig is cool stuff.

OpenStudy (anonymous):

my brain is fried from last semester but my teacher texted and did say I got a B so ill take that. :)

OpenStudy (anonymous):

ok ttyl I hope someone comes and helps you

OpenStudy (anonymous):

thanks for helping @radar

OpenStudy (anonymous):

If I just had.. "Need help finding area" as my question title.. more people would come in.....

OpenStudy (anonymous):

@thomaster @Mertsj maybe you can help? I got 2/3 sq units as an answer. Thanks in advance!

OpenStudy (loser66):

y = -x^2 +1, is it not a downward parabola with the vertex is (0,1)? so, the area under the curve is calculated by int from -1 to 1 of this curve?

OpenStudy (radar):

Sorry I couldn't help, I keep telling myself to review calculus, but never seem to get to it.

OpenStudy (anonymous):

so I need to integrate from -1 to 1?

OpenStudy (anonymous):

ahh since this is an ellipse?

OpenStudy (loser66):

you have parametric equations, and bink give you the equation above, To me, it's good enough to calculate. and Yes, take integral from -1 to 1 of -x^2 +1 parabola, not ellipse

OpenStudy (anonymous):

the answer says find the area enclosed by the ellipse.. is why i said that

OpenStudy (anonymous):

I am trying to use the formula A=int (a,b) y(dx/dt) dt

OpenStudy (anonymous):

do I not just start by calculating dx/dt?

OpenStudy (anonymous):

When I graph this parametric curve.. it does not cross the y axis?

OpenStudy (loser66):

I don't see any reason to go to multivariate problem. y = -x^2 +1, clear enough to have a curve,; clear enough to take integral

OpenStudy (loser66):

why not? y= 1- t^6, t =0 , y = 1 x = t^3 , t=0 x =0 it perfectly matches with y = -x^2 +1 you see, x =0 , y =1 is vertex of parabola y = -x^2 +1

OpenStudy (anonymous):

so the area is 4/3?

OpenStudy (anonymous):

I just plugged it into my graphing utility as a parametric curve.. and it only graphed in for positive values of x. I reset the domain.... and it still would not graph negative values of x.

OpenStudy (loser66):

I don't know, I don't calculate it, just give you my opinion

OpenStudy (anonymous):

I really need to use multivariate procedures on this since that is my class.

OpenStudy (anonymous):

oops not an ellipse, you are right.. just a curve.

OpenStudy (anonymous):

@Loser66 thank you for your help! I understand the limits of integration now... the problem worked out the same whether I used the multivariate method or the method you showed. Sorry, I was reading 2 problems at once.

OpenStudy (loser66):

how genius you are. I read just 1problem/time. And when I got problem, I want to cry with this only. YOU, 2problems/ time. ha!!!! jealous

OpenStudy (anonymous):

only 2 problems because I had to continue with my HW until help came.. which was a long time. It seems like the higher I go in math the more time it takes here to get help. Now trying to tackle arc length... long problems. Thanks again for your help!

OpenStudy (loser66):

hihihi, gooooooodddd luck.